OpenStudy (anonymous):

How many solutions does the system have? x = -4y + 4 2x + 8y = 8 A. one solution B. two solutions C. infinitely many solutions D. no solution

OpenStudy (akashdeepdeb):

How many do you think it does?

OpenStudy (akashdeepdeb):

@MonsterDoll ? :)

OpenStudy (nurali):

x = -4y + 4--(1) 2x + 8y = 8--(2) put x in equation (2) 2(-4y+4)+8y=8 -8y+8+8y-8=0 0=0

OpenStudy (akashdeepdeb):

x = -4y + 4 2x + 8y = 8 If you see the second equation is just the first equation multiplied by 2! So if there are lines co-inciding they'd have INFINITE SOLUTIONS!! :D